Find the coordinates of the midpoint of a segment with each pair of endpoints.
I(0, -2), J(3/4, 5)

Answer by stanbon(48568)

Find the coordinates of the midpoint of a segment with each pair of endpoints.
I(0, -2), J(3/4, 5)
The x coordinate is (0+(3/4))/2 = (3/8)
The y coordinate is (-2+5)/2=(3/2)
The midpoint is (3/8,3/2)
Cheers,
Stan H.